OpenStudy (anonymous):

Find all solutions of the following linear congruence. \(3x\equiv2\mod7\) First of all, we notice that \((3,7)=1\). Therefore, we will only have \(1\) solution. We now need to obtain a solution of the linear, diophantine equation \(3x-7y=2\). The Euclidean algorithm gives: \(7=3\cdot2+1\), \(3=1\cdot3+0\). Hence, \(7\cdot1-3\cdot2=1\) and \(7\cdot2-3\cdot4=2\). Therefore, a particular solution to the diophantine equation is \(x_0=-4\), \(y_0=-2\) and all solutions of the linear congruences are given by \(x\equiv-4\equiv3\mod7\). Am I right?

myininaya (myininaya):

I have a way but you might now like it

myininaya (myininaya):

not*

OpenStudy (anonymous):

i am sure this is right, but i would write the following \[3x\equiv2\mod7\] \[3x\equiv9\mod7\] \[x\equiv3\mod7\]

myininaya (myininaya):

\[3x \equiv 2 \mod 7 =>3x=2+7k\] k=0, there is no integer x k=1, x=3 k=2, there is no integer x k=3, there is no integer x k=4, x=10 k=5, there is no integer x k=6, there is no integer x k=7, x=17 so on... so we have x=...,3,10,17,17+7,(17+7)+7, and so on...

OpenStudy (anonymous):

that is the usual gimmick. keep adding until you can divide to solve the congruence

myininaya (myininaya):

x=3+7i where i is an integer

myininaya (myininaya):

which is what satellite said

OpenStudy (anonymous):

in two lines lol

myininaya (myininaya):

shhh... don't make fun of my inefficient way ;)

OpenStudy (anonymous):

speed kills